Using your result in problem (30), and the fact th
Answer #1

Problem # 36

DATA:

\omega =500\, rev/min

t =3.5\, s

\omega _{0} =0

\theta =\,?

SOLUTION:

First, lets convert the angular speed:

500\, \frac{rev}{min} .\frac{1\, min}{60\, s}=\frac{25}{3}\, rev/s

The angular speed \omega is defined as:

\omega=\omega _{0}+\alpha t

Where:

\alpha : angular acceleration

t : time

\omega _{0} : Initial angular speed

\omega : angular speed

Isolating the angular acceleration :

\alpha =\frac{\omega -\omega _{0}}{t}

Replacing values given in DATA:

\alpha =\frac{25/3\, rev/s -0}{3.5\, s}

Therefore:

\alpha =2.38\, rev/s^{2}

Now, the number of rotations can be calculate as follows:

\theta =\theta _{0}+\omega _{0}t+\frac{\alpha t^{2}}{2}

Repacing values:

\theta =0+0\times (3.5\, s)+\frac{2.38\, rev/s^{2}\times (3.5\, s)^{2}}{2}

Therefore:

{\color{Blue} \theta =14.58\approx 15}

OPTION C